40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 4 00:10:43 2020, gautam, Update, General, Vent activities today IMG_0184.jpgIMG_0185.jpgIMG_0187.jpgIMG_0207.jpegIMG_0210.jpg
    Reply  Thu Nov 5 22:27:34 2020, gautam, Update, General, Vent activities today 
       Reply  Fri Nov 6 11:36:37 2020, gautam, Update, General, ETMY suspension eigenmodes ETMY_pkFitNaive.pdfETMY.tar.bz2
          Reply  Tue Nov 10 15:13:41 2020, rana, Update, General, ETMY suspension eigenmodes 
             Reply  Thu Nov 12 14:26:35 2020, gautam, Update, General, ETMY suspension eigenmodes ETMY.tar.bz2
       Reply  Sun Nov 8 20:22:35 2020, gautam, Update, General, Vent activities today IFOalignment.png
          Reply  Sun Nov 8 23:24:29 2020, Koji, Update, General, Vent activities today 
          Reply  Mon Nov 9 17:40:17 2020, gautam, Update, General, Pumpdown IMG_8937.jpgpumpdown.png
             Reply  Tue Nov 10 11:31:13 2020, Koji, Update, General, Pumpdown Screen_Shot_2020-11-10_at_11.30.21.png
Message ID: 15666     Entry time: Mon Nov 9 17:40:17 2020     In reply to: 15664     Reply to this: 15667
Author: gautam 
Type: Update 
Category: General 
Subject: Pumpdown 

[koji, rana, gautam]

1100 - EY chamber inspected, no issues were found --> EY heavy door on

1200 - OMC chamber was inspected. OM6 was marginally tweaked to bring the beam down a little in pitch, and also a little northwards in Yaw. --> Heavy door on.

1230 - Pumpdown started. Initially, the annuli volume was pumped down. The procedure calls for doing this with the small turbopumps. However, V7 was left open, and hence, in the process, the TP1 foreline pressure (=P2) hit ~30 torr. This caused TP1 to shutdown. We were able to restart it without issue. This case was not caught by the interlock code, which was running at the time. It should be recitified.

1330 - OMC breadboard clean optics and DCPD hardware were wrapped up and packed into tupperware boxes and stored along the south arm. OMC cavity itself, the OMMT, and the breadboard the OMC was sitting on are wrapped in foil/Ameristat and stored in cabinet S13, lower 2 shelves.

1915 - P1a = 0.5 torr pressure reached. Switched over to pumping the main volume with TP1, backed by TP2 and TP3, which themselves are backed by their respective dry pumps and also the AUX drypump for some extra oomph. All cooling fans available in the area were turned on and directed at the turbo pumps. RV2 was used to throttle the flow suitably.

It was at this point that we hit a snag - RV2 has gotten stuck in a partially open position, see Attachment #1. We can see that the thread doesn't move in response to turning the rotary dial. Fortunately, the valve is partially open, so the main volume continues to be pumped - see Attachment #2 for the full history of today's pumping. We are leaving the main volume pumped in this configuration overnight (TP1 pumping main volume backed by TPs 2 and 3, which are in turn backed by their respective drypumps and also the AUX dry pump). I think there is little to no risk of any damage to the turbo pumps, the interlocks should catch any anomalies. The roughing pumps RP1 and RP3 were turned off and that line was disconnected and capped.

What are our options?

  1. The main volume is able to reach the "nominal" pressure of 1e-5 torr, just takes longer.
  2. At some point, we may be able to pump the main volume directly with TP2 and TP3 - it is unclear at this point whether it's better to have the conductance limited TP1, which has a higher pumping capacity, or have the smaller TPs 2 and 3 pump through a larger conductance. 
  3. Depending on how low the ultimate pressure gets, we may be able to run the usual IFO activities until the replacement pressure gauges arrive in ~1 week, at which point we can vent the pumpspool (leaving the main volume isolated) and either repair this valve or replace it with one of the spares we have.

We need some vacuum experts to comment. Why did this happen? Is this an acceptable failure mode of the valve?

KA Ed:
2230 - P1a = 0.025 torr. The pressure is coming down with log-linear scale. x0.1 per 2.5 hours or so.

Attachment 1: IMG_8937.jpg  2.823 MB  Uploaded Mon Nov 9 20:46:36 2020  | Hide | Hide all
IMG_8937.jpg
Attachment 2: pumpdown.png  44 kB  Uploaded Mon Nov 9 20:59:57 2020  | Hide | Hide all
pumpdown.png
ELOG V3.1.3-